OpenStudy (anonymous):

Simplify the product using the distributive property. (-4h +2)(3h +7)

OpenStudy (anonymous):

The distributive property in this case implies "foiling" which is an acronym for "first-outside-inside-last"

OpenStudy (anonymous):

I personally find that confusing so I think, "each term in the first brackets needs to be multiplied by each term in the second brackets"

OpenStudy (anonymous):

(-4h+2)(3h+7)= -4h(3h)+-4h(7)+2(3h)+2(7)= -12h^2+-28h+6h+14= -12h^2-22h+14

OpenStudy (anonymous):

A. -12h2 - 22h + 14 B. -12h2 - 34h - 14 C. -12h2 + 34h - 14 D. -12h2 + 22h + 14 I got B? thats not right?

OpenStudy (anonymous):

I got A, but I may have made a mistake. I've explained my steps, see if you agree with my actions.

OpenStudy (theviper):

(-4h +2)(3h +7) -4h(3h+7)+2(3h+7) -12h^2-28h+6h+14

OpenStudy (anonymous):

\[(-4h+2)(3h+7)\]\[-4h(3h+7)+2(3h+7)\]\[-12h ^{2}-28h+6h+14\]\[-12h ^{2}-22h+14\]